Find the mass and center of mass of the lamina that occupies the region bounded by y=x+ 2 and y=x^2, where the density at any point is given by σ(x,y) = 3x^2.
1
Expert's answer
2020-11-05T16:39:36-0500

The intersection between the curves "y=x+2" and "y=x^2" is,


"x+2=x^2"


"x^2-x-2=0"


"x^2-2x+x-2=0"


"x(x-2)+1(x-2)=0"


"(x+1)(x-2)=0"


"x=-1,2"


So, the point of intersection between the curves are "(-1,1),(2,4)"


The sketch of the region of lamina is as shown in the figure below:





The mass of the lamina is evaluated as,


"m=\iint_{D}\sigma(x,y)dA"


"=\iint_{D}3x^2dA"


"=\int_{-1}^{2}\int_{x^2}^{x+2}3x^2dydx"


"=\int_{-1}^{2}3x^2(x+2-x^2)dx"


"=\int_{-1}^{2}(3x^3+6x^2-3x^4)dx"


"=[3(\frac{x^4}{4})+6(\frac{x^3}{3})-3(\frac{x^5}{5})]_{-1}^{2}"


"=\frac{45}{4}+2(9)-\frac{99}{5}"


"=\frac{189}{20}"


Therefore, the mass of the lamina is "m=\frac{189}{20}"


The "x" -coordinates of the center of mass is evaluated as,


"x_{bar}=\frac{1}{m}\iint_{D}x\sigma(x,y)dA"


"=\frac{20}{189}\iint_{D}x(3x^2)dA"


"=\frac{20}{189}\int_{-1}^{2}\int_{x^2}^{x+2}3x^3dydx"


"=\frac{20}{189}\int_{-1}^{2}3x^3(x+2-x^2)dx"


"=\frac{20}{189}\int_{-1}^{2}(3x^4+6x^3-3x^5)dx"


"=\frac{20}{189}[3(\frac{x^5}{5})+6(\frac{x^4}{4})-3(\frac{x^6}{6})]_{-1}^{2}"


"=\frac{20}{189}(\frac{99}{5}+\frac{45}{2}-\frac{63}{2})"


"=\frac{20}{189}(\frac{54}{5})"


"=\frac{8}{7}"


The "y" -coordinate of the center of mass is evaluated as,


"y_{bar}=\frac{1}{m}\iint_{D}y\sigma(x,y)dA"


"=\frac{20}{189}\iint_{D}y(3x^2)dA"


"=\frac{20}{189}\int_{-1}^{2}\int_{x^2}^{x+2}3x^2ydydx"


"=\frac{20}{189}\int_{-1}^{2}3x^2\frac{1}{2}[((x+2)^2-x^4)]dx"


"=\frac{20}{189}(\frac{3}{2})\int_{-1}^{2}x^2(x^2+4x+4-x^4)dx"


"=\frac{20}{189}(\frac{3}{2})\int_{-1}^{2}(x^4+4x^3+4x^2-x^6)dx"


"=\frac{20}{189}(\frac{3}{2})[\frac{x^5}{5}+4(\frac{x^4}{4})+4(\frac{x^3}{3})-\frac{x^7}{7}]_{-1}^{2}"


"=\frac{20}{189}(\frac{3}{2})[\frac{33}{5}+15+12-\frac{129}{7}]"


"=\frac{20}{189}(\frac{3}{2})(\frac{531}{35})"


"=\frac{118}{49}"


Therefore, the "y" -coordinate of the center of mass is "y_{bar}=\frac{118}{49}"



So, the coordinates of the center of mass is "(x_{bar},y_{bar})=(\frac{8}{7},\frac{118}{49})"